# GRIL-Calib: Targetless Ground Robot IMU-LiDAR Extrinsic Calibration Method using Ground Plane Motion Constraints

### v1

@@ Abstract
Targetless IMU-LiDAR extrinsic calibration methods are gaining significant attention as the importance of the IMU-LiDAR fusion system increases. Notably, existing calibration methods derive calibration parameters under the assumption that the methods require full motion in all axes. When IMU and LiDAR are mounted on a ground robot the motion of which is restricted to planar motion, existing calibration methods are likely to exhibit degraded performance. To address this issue, we present GRIL-Calib: a novel targetless Ground Robot IMU-LiDAR Calibration method. Our proposed method leverages ground information to compensate for the lack of unrestricted full motion. First, we propose LiDAR Odometry (LO) using ground plane residuals to enhance calibration accuracy. Second, we propose the Ground Plane Motion (GPM) constraint and incorporate it into the optimization for calibration, enabling the determination of full 6-DoF extrinsic parameters, including theoretically unobservable direction. Finally, unlike baseline methods, we formulate the calibration not as sequential two optimizations but as a single optimization (SO) problem, solving all calibration parameters simultaneously and improving accuracy. We validate our GRIL-Calib by applying it to three public real-world datasets and comparing its performance with that of existing state-of-the-art methods in terms of accuracy and robustness. Our code is available at https://github.com/Taeyoung96/GRIL-Calib.
@@
Calibration and identification, Sensor Fusion, Optimization, Mobile Robot.
@@ I INTRODUCTION
LiDAR (Light Detection And Ranging) acquires more accurate 3D information than other sensors, making it broadly applicable to a variety of robotic applications such as ego-motion estimation [1], localization [2] and mapping [3] system. However, LiDAR has the drawback that when it is applied to fast motion, it is vulnerable to motion distortion. Recently, to overcome the drawback, LiDAR has been used in conjunction with IMU (Inertial Measurement Unit) [4], [5], [6]. The IMU effectively compensates for the motion distortion of the LiDAR, as it measures acceleration and angular velocity at a high frequency. Obviously, accurate extrinsic calibration is a fundamental requirement in the IMU-LiDAR fusion system. Extrinsic calibration refers to the process of establishing the spatial transformations between IMU and LiDAR, thereby aligning the measurements from the two sensors. If the extrinsic calibration is not accurate enough, the fusion of IMU and LiDAR becomes less effective, leading to significant degradation in overall performance [7].
@@
  • We validate our GRIL-Calib on M2DGR [14], Hilti [15], and S3E [16] datasets where it achieves excellent results compared to state-of-the-art targetless methods. We have open-sourced [^1] https://github.com/Taeyoung96/GRIL-Calib our algorithm to benefit the broader robotics community.
@@ II Related Work
In early IMU-LiDAR calibration research, an arbitrary target is required for calibration [8], [17]. However, the target-based method is not practical, so targetless calibration methods have been studied more recently. Subodh et al. [18] formulate a motion-based calibration based on the Extended Kalman Filter (EKF). Additionally, Lv et al. [19] utilizes a continuous-time trajectory formulation based on B-Spline. Also, OA-LICalib [11] which extends their previous works [19], proposes two observability-aware modules designed to address the degenerate case. Unfortunately, the proposed observability-aware module is designed to only reject unnecessary motion. FAST-LIO2 [4] is an accurate LiDAR-inertial odometry (LIO) algorithm based on an Iterated error state Kalman filter (IESKF) which include extrinsic calibration parameters. Due to the strong nonlinearity, the performance of the LIO system is highly dependent on accurate initial states. To get an accurate estimate of the initial states, Zhu et al. [20] propose the optimization-based algorithm that aligns angular velocity and acceleration. Existing approaches [18], [11], [20] require sufficient motion in all axes of the sensor to obtain accurate results. These methods are practical when applied to handheld devices that can move freely in various directions. However, existing methods encounter challenges when applied to ground robots where motion is limited to planar motion. In contrast, our method utilizes ground observation that can easily be observed by a ground robot to obtain accurate 6-DoF extrinsic parameters, even in the absence of sufficient motion in all axes. Recently, an extended Kalman filter (EKF) based IMU-LiDAR calibration algorithm utilizing a GNSS observation model in planar motion has been proposed [7]. However, it requires an additional GNSS sensor, while our proposed algorithm calibrates IMU-LiDAR without additional sensors.
@@ III System Overview
  • Data Synchronization, aimed at determining the approximate time offset between IMU and LiDAR data using cross-correlation methods. Notably, this process aligns with the approach proposed in [20].
@@ IV-C IMU Processing and Synchronization with LiDAR Odometry
Similar to [20], we remove the noise $\boldsymbol{\varepsilon}_{g}$ and $\boldsymbol{\varepsilon}_{a}$ from the IMU measurements with a zero-phase filter [25] in 14. Also, we down-sample to get the IMU measurement at timestamp $k$ from the LiDAR.
@@
To achieve more precise synchronization between two sensors, we employ a coarse-to-fine approach to determine the time offset, denoted as $t^{I}_{L}$. As outlined in [20], $t^{I}_{L}$ is the sum of the coarse estimate $t_{c}$ and the fine estimate $t_{f}$, thereby defined as $t^{I}_{L}=t_{c}+t_{f}$. Similar to [20], we determine $t_{c}$ by using the cross-correlation method given by
@@ V-A Unified Temporal-Spatial Constraints
by compensating the time offset $k^{\prime}=k+t_{c}^{*}$ between two sensors, where $t_{c}^{*}$ is given in 16 and assuming the IMU angular acceleration $\boldsymbol{\bar{\alpha}}_{I_{k^{\prime}}}$ is constant over a small period of time $t_{f}$. This is described as follows, according to (19)
@@ V-C Single Optimization (SO) Process
and minimize the two residuals $\mathbf{r}_{w}$ and $\mathbf{r}_{a}$ along with our proposed residual $\mathbf{r}_{g}$ simultaneously, where $\Sigma_{\mathbf{r}_{w}},\Sigma_{\mathbf{r}_{a}},\Sigma_{\mathbf{r}_{g}}$ denotes the Mahalanobis distance for each covariance matrix. $\rho_{x}$, where $x=w,a,g$ refers to the scaled Cauchy kernel function assigned to each residual.
VI Experiments
VI-A Experiments Setup
@@
LI-Init [20] 1.120 0.383 1.305 0.311 2.108 0.236 0.904 0.263 0.784 0.223
GRIL-Calib (Proposed) 0.376 0.045 0.818 0.033 0.472 0.027 0.154 0.066 0.328 0.043
@@
Our implementation is based on C++, using ROS (Robot Operating System) on Ubuntu 18.04. We perform our experiments using a 4-core Intel® Core™ i5-6600 CPU and Titan XP GPU. We use Ceres-solver  [^2] http://ceres-solver.org/ to compute the non-linear optimization. We evaluated the algorithms using three public datasets: M2DGR [14], Hilti [15], and S3E [16], which cover both indoor and outdoor environments. Following the evaluation scheme used in [20], we only evaluate spatial extrinsic calibration parameters in terms of RMSE (Root Mean Squared Error), skipping the evaluation on IMU LiDAR time offset $t^{I}_{L}$. To show the effectiveness of our Gril-Calib, we compare our method with other recent IMU-LiDAR calibration methods which are named ILC [18], FAST-LIO2 [4], OA-LICalib [11], LI-Init [20]. All algorithms are reproduced in the same environment.
@@ VI-B Extrinsic Calibration Performance on the M2DGR
The calibration result on the M2DGR dataset is summarized in Table I. M2DGR uses Velodyne VLP-32C as the LiDAR and Handsfree A9 as the IMU. For a consistent comparison, the initial rotation and translation parameters are set to be the same for all algorithms. The initial rotation parameters are set to $\mathbf{R}^{I}_{L}=(-5.0,-5.0,5.0)$ with the unit in degrees. The initial translation parameters deviated by 0.402 meters from the ground truth based on RMSE, represented as $\mathbf{p}^{I}_{L}=(0.6,0.45,0.6)$. We initialized all remaining state vectors $\mathbf{t}_{f},\mathbf{b}_{g},\mathbf{b}_{a}$ to zeros. Data was collected from both indoor (Hall 04 and Lift 04) and outdoor (Gate 01, Rotation 02, and Street 08) environments, with partial visualizations provided in Fig 5. The calibration result on the M2DGR is summarized in Table I. The table shows that our Gril-Calib significantly outperforms the other competing methods, specifically reducing RMSE by 75% for rotations and 79% for translations compared to the baseline [20].
Figure 5: Trajectory visualization for Street 08 acquired outdoors (Left), and Hall 04 acquired indoors (Right). We evaluated the algorithm with data collected from a variety of environments.
@@
The extrinsic parameters of each calibration method are visualized in Fig. 6. Fig. 5(a) and Fig, 5(b) are the results of the rotation and translation matrices, respectively. In both figures, the dotted lines denote the ground truth of the extrinsic parameters. Thus, the closer to the dotted line, the better the calibration method is. ICL [18] is excluded from the figure because its error is significantly larger than that of the other methods.
VI-C Extrinsic Calibration Performance on the Others
S3E and HILTI data are used as additional datasets, and their results are summarized in Table II and Table III, respectively. Both of the datasets are captured indoors. The S3E dataset uses a Velodyne VLP-16 and Xsens MTi-30 IMU, while the HILTI dataset uses an Ouster OS0-64 LiDAR and AlphaSense IMU. As shown in Tables II and III, our algorithm outperforms other competing methods as in M2DGR, and it shows the reliable accuracy of our Gril-Calib across various robot platforms. We also observe that OA-LICalib fails in calibration due to a memory computation issue in the Baseline 3 sequence in Table III. The reason might be that OA-LICalib uses surfel maps for calibration and the computational cost rapidly increases with the number of surfels. In contrast, our algorithm works smoothly without any memory issues. The two tables along with Table I show the effectiveness and versatility of our approach in various real-world scenarios.
@@ VI-C Extrinsic Calibration Performance on the Others
ICL [18] 13.308 4.555 44.294 4.052
FAST-LIO2 [4] 1.964 0.196 1.892 0.188
OA-LICalib [11] $\times$ $\times$ 2.368 0.185
LI-Init [20] 0.374 0.092 2.584 0.194
GRIL-Calib (Proposed) 0.716 0.018 0.589 0.047
  • $\times$ means that the algorithm failed.
@@ VI-D Robustness to the Initial Guess
The extrinsic calibration usually requires initial parameter values, and it significantly affects the performance [27]. We conduct some experiments while varying initial guesses to see how robust the competing calibration methods are to the initial guesses. We use Gate 01 sequence in M2DGR. The results are shown in Fig 8. Fig 7(a) shows the change of the calibration error while fixing the initial $\mathbf{p}_{L}^{I}$ but changing initial $\mathbf{R}_{L}^{I}$. Fig 7(b) is the result when $\mathbf{R}_{L}^{I}$ is fixed but $\mathbf{p}_{L}^{I}$ is varied. Interestingly, OA-LICalib showed relatively good performance in the previous three datasets, but as the initial guess increases, its accuracy rapidly degrades. Instead, LI-Init demonstrates robust performance against the initial guess, but it is less accurate than our Gril-Calib due to the limited planar motion.
Figure 7: Evaluation of LiDAR-inertial odometry using different calibration algorithms.
Figure 8: Calibration performance with changes initial guess.
@@
(Proposed) 0.685 0.054
0.430 0.041
VI-E Application to IMU-LiDAR Fusion System
We compare the competing extrinsic calibration methods by applying the respective results to IMU-LiDAR fusion system and evaluating the corresponding odometry accuracy. In the experiment, we utilize Faster-LIO [6] as the IMU-LiDAR fusion system. The results in Table I used extrinsic parameters of each method. The experiment uses the hall 02 sequence since the sequence was not used in Subsection VI-B. The Absolute Pose Error (APE) is depicted from different calibration methods in Fig 7. We observe that the odometry obtained by our Gril-Calib outperforms those of other methods. In particular, our Gril-Calib achieves 8.6% improvement from LI-Init and 9.6% improvement from OA-LICalib. This experiment shows the importance of accurate calibration when applied to IMU-LiDAR fusion systems.
Figure 9: Comparative visualization of the z-component of a trajectory using Rotation 02 sequence in M2DGR.
VI-F Ablation Study
We ablate three key modules comparing the result with that of our baseline [20]. The results are summarized in Table IV. On the first row of the table, ‘GPM’, ‘SO’, and ‘LO w/ GP’ denote the three contributions made in our paper. That is, they imply ‘calibration using GPM constraints’, ‘calibration via single optimization’, and ‘LO with proposed ground plane residual’, respectively. Notably, SO observed a 34% decrease in RMSE for translation calibration. This reduction highlights that $\mathbf{R}^{I}_{L}$ and $\mathbf{p}^{I}_{L}$ are highly correlated and the separate optimization which ignores the correlation would lead to degraded accuracy. Our LO method with the proposed ground plane residual, effectively minimizes z-axis drift in the LO, as shown in Fig 9. This drift reduction leads to more stable and accurate measurements, thereby enhancing the overall precision of our calibration approach.

### v2

@@ Abstract
Targetless IMU-LiDAR extrinsic calibration methods are gaining significant attention as the importance of the IMU-LiDAR fusion system increases. Notably, existing calibration methods derive calibration parameters under the assumption that the methods require full motion in all axes. When IMU and LiDAR are mounted on a ground robot the motion of which is restricted to planar motion, existing calibration methods are likely to exhibit degraded performance. To address this issue, we present GRIL-Calib: a novel targetless Ground Robot IMU-LiDAR Calibration method. Our proposed method leverages ground information to compensate for the lack of unrestricted full motion. First, we propose LiDAR Odometry (LO) using ground plane residuals to enhance calibration accuracy. Second, we propose the Ground Plane Motion (GPM) constraint and incorporate it into the optimization for calibration, enabling the determination of full 6-DoF extrinsic parameters, including theoretically unobservable direction. Finally, unlike baseline methods, we formulate the calibration not as sequential two optimizations but as a single optimization (SO) problem, solving all calibration parameters simultaneously and improving accuracy. We validate our GRIL-Calib by applying it to various real-world datasets and comparing its performance with that of existing state-of-the-art methods in terms of accuracy and robustness. Our code is available at https://github.com/Taeyoung96/GRIL-Calib.
@@
Sensor Fusion, Calibration and Identification.
@@ I INTRODUCTION
LiDAR (Light Detection And Ranging) acquires more accurate 3D information than other sensors, making it broadly applicable to a variety of robotic applications such as ego-motion estimation [1], localization [2] and mapping [3] system. However, LiDAR has the drawback that when it is applied to fast motion, it is vulnerable to motion distortion. Recently, to overcome the drawback, LiDAR has been used in conjunction with IMU (Inertial Measurement Unit) [4], [5], [6]. The IMU effectively compensates for the motion distortion of the LiDAR, as it measures acceleration and angular velocity at a high frequency. Obviously, accurate extrinsic calibration is a fundamental requirement in the IMU-LiDAR fusion system. Extrinsic calibration refers to the process of establishing the spatial relationship, including both translation and rotation, between multiple sensors. This process ensures that measurements from different sensors are accurately interpreted using the same coordinate frames. If the extrinsic calibration is not accurate enough, the fusion of IMU and LiDAR becomes less effective, leading to significant degradation in overall performance [7].
@@
  • We validated GRIL-Calib on a variety of datasets, achieving significant results compared to state-of-the-art targetless methods. We have open-sourced [^1] https://github.com/Taeyoung96/GRIL-Calib our algorithm to benefit the broader robotics community.
@@ II Related Work
In early IMU-LiDAR calibration research, an arbitrary target is required for calibration [8], [14]. However, the target-based method is not practical, so targetless calibration methods have been studied more recently. Subodh et al. [15] formulate a motion-based calibration based on the Extended Kalman Filter (EKF). Additionally, Lv et al. [16] utilizes a continuous-time trajectory formulation based on B-Spline. Also, OA-LICalib [11] which extends their previous works [16], proposes two observability-aware modules designed to address the degenerate case. Unfortunately, the proposed observability-aware module is designed to only reject unnecessary motion. FAST-LIO2 [4] is an accurate LiDAR-inertial odometry (LIO) algorithm based on an Iterated error state Kalman filter (IESKF) which include extrinsic calibration parameters. Due to the strong nonlinearity, the performance of the LIO system is highly dependent on accurate initial states. To get an accurate estimate of the initial states, Zhu et al. [17] propose the optimization-based algorithm that aligns angular velocity and acceleration. Existing approaches [15], [11], [17] require sufficient motion in all axes of the sensor to obtain accurate results. These methods are practical when applied to handheld devices that can move freely in various directions. However, existing methods encounter challenges when applied to ground robots where motion is limited to planar motion. In contrast, our method utilizes ground observation that can easily be observed by a ground robot to obtain accurate 6-DoF extrinsic parameters, even in the absence of sufficient motion in all axes. Recently, there are also several approaches to calibrating different sensors on ground robots. Zuniga-Noel et al. [18] proposed a two-step multi-sensor calibration method for mobile robots using a camera and 2D laser. Also, an IMU-LiDAR calibration algorithm utilizing a GNSS sensor in planar motion has been proposed [7] and [19]. However, they require an additional GNSS sensor, while our proposed algorithm calibrates IMU-LiDAR without additional sensors.
@@ III System Overview
  • Data Synchronization, aimed at determining the approximate time offset between IMU and LiDAR odometry using cross-correlation methods that analyze their magnitude of angular velocity. Notably, this process aligns with the approach proposed in [17].
@@ IV-C IMU Processing and Synchronization with LiDAR Odometry
Similar to [17], we remove the noise $\boldsymbol{\varepsilon}_{g}$ and $\boldsymbol{\varepsilon}_{a}$ from the IMU measurements with a zero-phase filter [24] in (14). Also, we down-sample to get the IMU measurement at timestamp $k$ from the LiDAR.
@@
Since the IMU and LiDAR Odometry are not synchronized, we utilize the cross-correlation between their respective angular velocities to quantify the similarity between their magnitudes. To achieve more precise synchronization between two sensors, we employ a coarse-to-fine approach to determine the time offset, denoted as $t^{I}_{L}$. As outlined in [17], $t^{I}_{L}$ is the sum of the coarse estimate $t_{c}$ and the fine estimate $t_{f}$, thereby defined as $t^{I}_{L}=t_{c}+t_{f}$. Similar to [17], we determine $t_{c}$ by using the cross-correlation method given by
@@ V-A Unified Temporal-Spatial Constraints
by compensating the time offset $k^{\prime}=k+t_{c}^{*}$ between two sensors, where $t_{c}^{*}$ is given in (16) and assuming the IMU angular acceleration $\boldsymbol{\bar{\alpha}}_{I_{k^{\prime}}}$ is constant over a small period of time $t_{f}$. This is described as follows, according to (19)
@@ V-C Single Optimization (SO) Process
and minimize the two residuals $\mathbf{r}_{w}$ and $\mathbf{r}_{a}$ along with our proposed residual $\mathbf{r}_{g}$ simultaneously, where $\Sigma_{\mathbf{r}_{w}},\Sigma_{\mathbf{r}_{a}},\Sigma_{\mathbf{r}_{g}}$ denotes the Mahalanobis distance for each covariance matrix. $\rho_{x}$, where $x=w,a,g$ refers to the scaled Cauchy kernel function assigned to each residual. To determine the scale parameters for the Cauchy kernel function, we utilize a grid search in a predefined parameter space. We found the optimal parameter combination of $\rho_{w}:\rho_{a}:\rho_{g}=10:1:5$ that was used consistently across all experiments.
@@
LI-Init [17] 1.120 0.383 1.305 0.311 2.250 0.287 3.083 0.295 0.784 0.223
GRIL-Calib (Proposed) 0.501 0.039 0.878 0.021 0.559 0.034 0.873 0.085 0.639 0.125
@@
  • The following numbers under the sequence represent segments of data.
For data accumulation, we start collecting once the robot is ready after initializing the Madgwick filter. The end time for data collection is determined similarly to [17], ensuring the robot moves enough for accurate calibration. However, unlike [17], we only consider the largest singular value of the Jacobian matrix $\mathbf{J_{R}}$ of $\mathbf{R}^{I}_{L}$ since the ground robot is attached to the ground and moves only in planar motion.
VI Experiments
VI-A Experiments Setup
Our implementation is based on C++, using ROS on Ubuntu 18.04. We perform our experiments using a 4-core Intel® Core™ i5-6600 CPU and Titan XP GPU. We use Ceres-solver  [^2] http://ceres-solver.org/ to compute the non-linear optimization. We evaluated the algorithms using various datasets with flat ground. Following the evaluation scheme used in [17], we only evaluate spatial extrinsic calibration parameters in terms of RMSE (Root Mean Squared Error), skipping the evaluation on IMU-LiDAR time offset $t^{I}_{L}$. To show the effectiveness of our Gril-Calib, we compare our method with other recent IMU-LiDAR calibration methods which are named ILC [15], FAST-LIO2 [4], OA-LICalib [11], LI-Init [17]. All algorithms are reproduced in the same environment.
@@ VI-B Extrinsic Calibration Performance on the M2DGR
M2DGR uses Velodyne VLP-32C as the LiDAR and Handsfree A9 as the IMU. For the initial parameter settings, we decide to set values that are far enough away from the ground truth. The initial rotation parameters are set to $\mathbf{R}^{I}_{L}=(-5.0,-5.0,5.0)$ using Euler angle conversion, with units in degrees. The initial translation parameters deviated by 0.402 meters from the ground truth based on RMSE, represented as $\mathbf{p}^{I}_{L}=(0.6,0.45,0.6)$. We initialized all remaining state vectors $\mathbf{t}_{f},\mathbf{b}_{g},\mathbf{b}_{a}$ to zeros. Data was collected from both indoor (Hall 04 and Lift 04) and outdoor (Gate 01, Rotation 02, and Street 08) environments. The calibration result on the M2DGR is summarized in Table I. The table shows that our Gril-Calib significantly outperforms the other methods. The extrinsic parameters of each calibration method are visualized in Fig. 5. Fig. 5(a) and Fig, 5(b) are the results of the rotation and translation matrices, respectively. In both figures, the dotted lines denote the ground truth of the extrinsic parameters. Thus, the closer to the dotted line, the better the calibration method is. ILC [15] is excluded from the figures because its error is significantly larger than that of the other methods. Additionally, we report the computation time of the algorithm to evaluate its efficiency in the supplementary material.
@@
ILC [15] 12.496 1.622 44.294 4.052
FAST-LIO2 [4] 2.533 0.206 1.892 0.188
OA-LICalib [11] 3.979 0.448 2.368 0.185
LI-Init [17] 1.652 0.191 2.584 0.194
GRIL-Calib (Proposed) 0.069 0.018 0.589 0.047
VI-C Extrinsic Calibration Performance on the Others
S3E and HILTI data are used as additional real-world datasets, and their results are summarized in Table II and III. Both of the datasets are captured indoors with flat ground. The S3E dataset uses a Velodyne VLP-16 and Xsens MTi-30 IMU, while the HILTI dataset uses an Ouster OS0-64 LiDAR and AlphaSense IMU. Similar with Section VI-B, we set the initial value settings as follows. In HILTI, $\mathbf{R}^{I}_{L}=(-179.91,-1.0,-5.0)$ and $\mathbf{p}^{I}_{L}=(-0.1,-0.4,0.5)$. In S3E, $\mathbf{R}^{I}_{L}=(0.0,0.0,5.0)$ and $\mathbf{p}^{I}_{L}=(-0.1,-0.2,-0.4)$ respectively. As with other experiments, we initialized all remaining state vectors to zeros. As shown in Tables II and III, our algorithm outperforms other competing methods and it shows the reliable accuracy of our Gril-Calib across various scenarios. Due to the paper limitation, experiments on the custom dataset, tests conducted on rough terrains, and the application of extrinsic calibration to the IMU-LiDAR Fusion System, are provided in supplementary material.  [^3] https://github.com/Taeyoung96/GRIL-Calib
@@ VI-D Robustness to the Initial Guess
The extrinsic calibration usually requires initial parameter values, and it significantly affects the performance [29]. We conduct some experiments while varying initial guesses to see how robust the competing calibration methods are to the initial guesses. In our experiments, we set the initial values to be sufficiently far from the ground truth. We utilize three different sequences, Gate 01, Rotation 02, and Hall 04, and average them. We progressively set the initial values that are farther from the ground truth to visualize the results of different algorithms for changing the initial setting. The results are shown in Fig 6. Fig 6(a) shows the change of the calibration error while fixing the initial $\mathbf{p}_{L}^{I}$ but changing initial $\mathbf{R}_{L}^{I}$. Fig 6(b) is the result when $\mathbf{R}_{L}^{I}$ is fixed but $\mathbf{p}_{L}^{I}$ is varied. We observe that the accuracy of our GRIL-Calib is less affected by the choice of initial parameters.
@@
(Proposed) 0.985 0.216
1.203 0.213
0.695 0.067
1.109 0.073
0.914 0.204
0.690 0.061
  • $\cdot$ GPM: Add Ground Plane Motion Constraints
  • $\cdot$ SO: Single Optimization Process
  • $\cdot$ LO w/ GP: LiDAR Odometry with proposed Ground Plane Residual
Figure 6: Calibration performance with changes initial guess.
Figure 7: Comparative visualization of the z-component of a trajectory using Rotation 02, Hall 04, Gate 01 sequence in M2DGR. (Top to bottom)
VI-E Ablation Study
We ablate three key modules comparing the result with that of our baseline [17] using whole M2DGR sequences. The results are summarized in Table IV. Adding each module sequentially improved calibration performance. Specifically, the incorporation of ‘GPM’ resulted in improvements of 16.9% and 70% in the rotation and translation parameters, respectively, compared to the [17]. The GPM is particularly effective for the translation parameter in the unobservable direction, resulting in significant performance enhancements. Further improvements were observed when calibration is performed by adding Single Optimization (‘SO’), with increases of 51% in rotation and 23% in translation parameters. This improvement underscores the high correlation between $\mathbf{R}^{I}_{L}$ and $\mathbf{p}^{I}_{L}$, and highlights that separate optimization, which overlooks this correlation, compromises accuracy. Finally, adding the ground plane residual in LiDAR odometry (‘LO w/ GP’) further enhances performance. The reduction in LiDAR Odometry drift with the addition of ground plane residual is shown in Fig 7. The reduced drift in the odometry allows for more accurate LiDAR angular velocity and acceleration, which contributed to the improved accuracy of the calibration.